# Generalized Label-Efficient 3D Scene Parsing via Hierarchical Feature Aligned Pre-Training and Region-Aware Fine-tuning

### v1

@@ 3 Proposed Methodology
Figure 3: The fine-tuning paradigm of our proposed WS3D++, i.e. WS3D. WS3D [4] consists of three modules: 1. The unsupervised region-level energy-based optimization guided by boundary labels; 2. The unsupervised multi-stage region-level contrastive learning with high confidence; 3. The supervised region-level semantic contrastive learning with labeled data. The backbone network adopts encoder-decoder structures. The weights of the backbone network are shared in the supervised and unsupervised branches. Integrated with the proposed pre-training paradigm illustrated in Figure 2, by our proposed hierarchical feature aligned pre-training and regional fine-tuning, more effective label-efficient learning as well as open-vocabulary learning is realized.
@@ 3.1 Hierarchical Vision-Language Knowledge Associations and Distillations for Pre-training
Multi-view rendering. To obtain paired 2D-3D representations, we propose leveraging multi-view rendering to obtain paired 2D views from 3D point cloud scenes. The pairing process consists of two steps, the first is to convert point cloud scenes into meshes and the second is to render 2D images based on the different views of the 3D meshes. In terms of point-to-mesh transformation, we utilize the Delaunay triangulation approach [107, 108] to convert the point cloud into meshes, which is demonstrated as a very effective method for surface reconstruction. It connects the points in point cloud scenes by forming triangles that satisfy the Delaunay criterion which guarantees no point lies inside the circumcircle of any triangle. This method generates a triangle mesh that approximates the surface of the point cloud [109]. In terms of mesh-to-image transformation, we leverage the rendering pipeline including the vertex transformation, projection, and rasterization, as well as shading. We directly use the rendering library OpenGL [110] to render images from the meshes. The process involves projecting the 3D vertices onto a 2D image plane based on camera parameters and applying shading and lighting calculations of 3D meshes to determine the specific color of each pixel. By this simple design, the world-to-camera extrinsic matrix $T_{e}$ containing both rotation and translation information between the 2D pixels and 3D points can be easily obtained.
@@
United 2D and 3D Proposal Generation. As shown in Fig. 6, according to our experiments, we found that it is difficult to directly find the object-level information merely based on the 2D proposals due to the great information loss while conducting rendering, the proposals provided by the 2D region proposal network (RPN) [112] can not effectively capture the 3D object information within the scene. On the other hand, merely relying on the 3D proposal provided by the 3D RPN [113] still can not guarantee accurate proposal generation for the fact that some objects are too adjacent in their geometry. To this end, we propose to leverage the union of 2D and 3D RPN to capture intact and all-inclusive proposals within the 3D scenes. Denote the region proposals as $R_{2D}$ and $R_{3D}$ respectively, the final holistic proposal generation $R_{H}$ is formulated as $R_{H}=R_{2D}\cup R_{3D}$. According to our experiments, this simple design can considerably boost the performance for the fact that the objects are more clustered and closely distributed within the indoor scenes. Taking the union of 2D and 3D object proposals into consideration also guarantees that the optimization merely considers the regions where the object really exists and prevents the models from taking the pure background into consideration during the optimizations.
Scene-level Sentence-3D Matching. We perform image-language alignment at both the object level and the scene level to pre-train the backbone network, which captures very accurate language-2D association for rendered images. The alignment at the object level allows the capturing of the fine-grained object-to-semantic concept matching, while the alignment at the scene level captures both holistic scene-level information and the dependent relations between the objects and holistic 3D scenes. For example, the bed exists in the bedroom while the sofa and TV lie in the living room with a clear owner-member relationship. To capture the object proposal information from the rendered images, we utilize the region proposal network (RPN) [112] to obtain the class-agnostic regions. For the text descriptor within the region, we use three most similar noun phase extract by it given by CLIP [9], which captures the scene level and the object level image patches with the most relevant texture description. Then we pass through both the original scene-level image as well as the obtained region-level image into the frozen CLIP visual encoder to obtain the scene-level visual feature $\textbf{I}_{\textbf{g}}$ as well as the object-level visual feature $\textbf{I}_{\textbf{l}}$, respectively. As shown in Figure 2, in order not to influence the well-learned feature representations provided by the CLIP encoder, we freeze the backbone and merely add a linear projection layer at the end of the network to facilitate the adaptive hierarchical feature-aligned pre-training.
Global and Regional Word-3D Matching. At the next step, we perform regional word-to-3D matching. We use the CLIP image encoder to generate a set of local feature embeddings $\textbf{F}^{l}_{3D}$ for the rendered proposal-level image patches, and the set of global embedding $\textbf{F}^{g}_{3D}$ for the whole scene. While the text encoder extracts the local textual embeddings $\textbf{F}^{l}_{T}\in\mathbb{R}^{H\times D}$ and the global textual embeddings $\textbf{F}^{g}_{T}$. The operation can be interpreted as a bi-directional operation, which means that for each region proposal image patch, we find the textual concept that matches best with the semantics of the region. And for each textual concept, we find the multiple regions that correspond with it. Next, we calculate the cosine similarity $\mathcal{G}_{sim}$ between the 3D visually encoded features and the textual features:
$\textbf{S}(\textbf{I},\textbf{T})=\mathcal{G}_{sim}(\textbf{F}^{l}_{3D},\textbf{F}^{l}_{T})+\mathcal{G}_{sim}(\textbf{F}^{g}_{3D},\textbf{F}^{g}_{T})$ (1)
@@
Contrastive Optimizations. Note that when conducting contrastive learning, we regard textual features as anchors because textual descriptions are highly semantic and contain rich information, whereas the images contain too much low-level information and pixel-level details. Denote the $\textit{{F}}^{\;b,+}_{3D}$ and the $\textit{{F}}^{\;b,-}_{3D}$ as the positive and the negative features with respect to the anchor textual feature $\bm{F}_{T}^{a}$, respectively, the designed contrastive loss is formulated as follows:
@@
$\mathcal{L}^{total}_{Ctr}=\mathcal{L}^{G}_{Ctr}+\mathcal{L}^{L}_{Ctr}$ (3)
3D Language Distillation. We utilize the $\mathcal{KL}$ divergence as the distillation loss to further distill the knowledge from 2D feature space to 3D. Compared with the mean square error loss such as the $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ losses, the $\mathcal{KL}$ divergence has improved regression capacity and ensured smoother gradient [115], which to some extent overcomes overfitting problems while distilling knowledge. The distillation loss $\mathcal{L}^{Dist}_{\mathcal{KL}}$ is formulated as:
@@
According to our extensive experiments, our simple pre-training approach provides well-aligned vision-language-3D co-embedding. It largely facilitates vision-language-associated knowledge transfers from 2D to 3D, boosting both the label efficiency and the final recognition capacity of novel categories.
@@ 3.2 Region-Aware Fine-tuning
During the fine-tuning stage, our proposed framework consists of three subparts for the network optimization: 1. Unsupervised energy-based loss guided by boundary awareness and highly confident network predictions for unlabeled data, which is discussed in 3.2.1; 2. Unsupervised multi-stage region-level contrastive learning with highly confident predictions for unlabeled data, which is discussed in 3.2.2. 3. Supervised semantic contrastive learning for labeled data, which is discussed in 3.2.3. The three modules above are integrated jointly into the optimization function for network training to accomplish the final downstream detection or segmentation tasks with a very limited labeled data and a large amount of unlabeled data.
3.2.1 Unsupervised Region-level Boundary Awareness
Energy-function-based conditional random field segmentation is proposed in [29] and has been widely applied. However, it works in a fully supervised manner and does not explicitly consider the semantic boundary information, which is a great indication of semantic partitions in 3D scenes. In this Subsection, we develop a boundary-aware energy-based loss for unsupervised learning. As shown in Fig. 1, to obtain robust boundaries for unlabeled 3D points, we first perform 3D over-segmentation [34] and also extract boundary points using an off-the-shelf semantic boundary prediction network, which is both subsequently used as the conditions to define the boundary regions for 3D points. Then, we propose a region-level energy-based loss based on obtained boundary region labels.
Point Cloud Over-Segmentation. To obtain boundary regions, and facilitate subsequent region-level affinity computation and region-level contrastive learning, we first perform a region-level coarse clustering-based on point cloud over-segmentation. The previous method depended on the region growing [34] for over-segmentation, which relied heavily on the accurate normal estimation and was easily influenced by noises. In our work, we choose to use normal, curvature, and point feature histogram (PFH) [116] simultaneously to provide initial over-segmentation. The detailed procedures are given in the Appendix. And the over-segmentation results are shown in Fig. 1 and 5. Denote original point clouds as $P_{in}$. After over-segmentation, they are partitioned into $Q$ subregions $S=\{s_{1},s_{2},...,s_{q}\}$, where ${s_{i}}\cap{s_{k}}=\emptyset$ for any ${s_{i}}\neq{s_{k}}$ as shown in Fig. 1 and 5.
Boundary Points Extraction. As shown in Fig. 1, in addition to the over-segmentation result, we first extract the semantic boundary points to further identify boundary regions. The semantic boundary often indicates the distinguishment between various semantic classes. We extract semantic boundary points by JSENet [117], as shown in Fig. 3. Note that the JSENet is trained with diverse ratios of labeled data in weakly supervised manners. As for training, we first define semantic boundary points from the limited labeled scenes as ground truth. With the definition of the ground truth boundary points, we design the loss following JSENet except substituting the binary cross entropy loss $L_{bce}$ with the focal loss $L_{foc}$ [118] to tackle large class imbalances between the boundary points and non-boundary points. $L_{foc}$ can be formulated as follows:
$L_{foc}=-\frac{1}{N_{i}}\sum_{i=1}^{N_{i}}{(1-b^{gt}_{i})}^{\alpha}log(b_{i})+{(b^{gt}_{i})^{\alpha}}log(1-b_{i}),\vspace{-0.58 mm}$ (6)
where $b_{i}$ denotes the binary predicted boundary map and $b^{gt}_{i}$ denotes the ground truth boundary map. $N_{i}$ is the total number of input points for training. We select $\alpha$=2 based on the original design [118]. After its convergence, we apply the trained network to the remaining unlabeled scenes to obtain their boundary points. Examples of predicted boundary points of ScanNet [20] are shown in Fig. 5, which clearly reveal distinctions between diverse semantic classes.
Figure 4: Qualitative semantic segmentation results of the proposed WS3D++ on SemanticKITTI validation set the with the 5% labeling percentage, compared with the fully supervised state-of-the-art Cylinder3D [119], and BAAF-Net [89] with the semantics indicated by different colors. The red circles highlight the final performance difference between diverse approaches.
Boundary Labels. After extracting semantic boundary points, we utilize them as labels of discrimination between diverse semantic categories. As shown in Fig. 1, denote the $j_{th}$ adjacent regions of the center region $s_{i}$ as $s_{i,j}$. The adjacent region query is realized by fast Octree-based K-nearest neighbor search [120]. Then, we determine the two adjacent regions as boundary regions if both $s_{i}$ and $s_{i,j}$ contain boundary points. The label for boundary region $w_{i,j}$ is designed as:
$w_{i,j}=\begin{cases}1&\textit{if $s_{i}$, $s_{i,j}$ both contain boundary points;}\\ 0&\textit{otherwise.}\end{cases}$ (7)
The label $w_{i,j}$ denotes semantic boundaries of adjacent regions, which is then used to guide the optimization of the energy function for the downstream segmentation task.
Energy Loss Guided by Boundary Labels. As shown in Fig. 3, we first perform data augmentation (detailed in the Appendix) for the input point clouds $\textbf{P}_{in}$ to obtain two transformed point clouds $\textbf{P}_{in,1}$ and $\textbf{P}_{in,2}\in\mathbb{R}^{N\times C_{in}}$, where N is the numbers of points. $C_{in}$ and $C_{out}$ are the numbers of input and output feature channels, respectively. Utilizing the backbone network SparseConv [59], we can obtain point cloud predictions $\textbf{P}_{out,1}$ and $\textbf{P}_{out,2}$. Then applying region-level max pooling on the same subregions, we obtain the predicted classes $\textbf{P}_{1}$, and $\textbf{P}_{2}\in\mathbb{R}^{M\times C_{out}}$ of the specific subregions. $\textbf{P}=\{p(s_{1}),p(s_{2}),...p(s_{i}),...,p(s_{R})\}$, where $R$ is the total number of regions obtained by over-segmentation. Denote the prediction of the $j_{th}$ neighboring region of the center region $s_{i}$ as $p(s_{i,j})$. Taking the unary network prediction and pairwise affinity between the neighbouring region into account, inspired by conditional random field (CRF) in DeepLab [29], we formulate the optimization energy function $E_{sum}$ as follows for label-efficient learning:
$E_{sum}=\sum_{i}E_{i}(s_{i})+\sum_{i<j}^{adjacent}E_{i,j}(s_{i},s_{i,j}).$ (8)
The first unary network prediction item $E_{i}(s_{i})$ is the entropy regularization term. It encourages region-level prediction with high confidence, which also facilitates the contrastive learning introduced subsequently, which is formulated as:
$E_{i}(s_{i})=-log\,p(s_{i}).$ (9)
We propose the pairwise affinity term of $E_{sum}$ as:
$\begin{split}E_{i,j}(s_{i},s_{i,j})=H_{i,j}w_{i,j}[\epsilon-\|p(s_{i})-p(s_{i,j})\|]^{2}_{+}\\ +H_{i,j}(1-w_{i,j})\|p(s_{i})-p(s_{i,j})\|^{2}.\end{split}$ (10)
$H_{i,j}$ is designed as the confidence indicator. $H_{i,j}=1$, if the probabilities which produce $p(s_{i})$ and $p(s_{i,j})$ are both larger than a threshold $\gamma$. Otherwise, it equals 00. $\epsilon$ can be any value in the range of $(0,1)$. $p(s_{i})$ and $p(s_{i,j})$ are the semantic predictions for the $i_{th}$ center region and the $j_{th}$ neighbouring region, respectively. And $[x]_{+}$ is the maximum function $max(0,x)$. For adjacent boundary regions, we encourage their confident semantic or instance predictions to be different (i.e. larger $\|p(s_{i})-p(s_{i,j})\|$); while for the non-boundary adjacent regions, we force their semantic predictions to be the same (i.e. $\|p(s_{i})-p(s_{i,j})\|=0$). Different from the traditional energy function in DeepLab[29], which used handcrafted features to compare similarities, we propose to use the learned boundary region labels to guide the network’s confident max-pooled region-level predictions. Therefore, the proposed boundary-aware energy function better encourages semantic separations at boundaries. Furthermore, we only consider pairwise affinity between adjacent regions instead of all pixel pairs, which greatly reduces computation costs, and avoids noises induced by distant unrelated pairs in the meanwhile.
Figure 5: Qualitative instance segmentation results of proposed WS3D on ScanNet [20] with fully supervised state-of-the-art and the previous state-of-the-art GPC [121] in the 5% labeling ratio, compared with fully supervised arts. Different instances are indicated by different colors. And the intermediate over-segmentation results for obtaining regions and boundary predictions, with the boundary points indicated by the color blue.
3.2.2 Region-level Instance Discrimination
After applying the entropy regularization term $E_{i}(s_{i})$, we can obtain predictions at the region level with high confidence. Note that confident region-level predictions further improve the latent feature discrimination capacity of the network, which makes contrastive learning in the latent space feasible. Therefore, we further propose a multi-stage region-level contrastive learning for unlabeled data. Compared to previous work using only contrastive learning with low-level geometric registrations [27], our work unleashes potentials of contrastive learning with instance discrimination to enhance versatile feature learning in latent space.
The key to successful semantic/instance segmentation is maintaining discriminative feature representations at different stages of the backbone network [122]. Inspired by the pyramidal feature network [122, 123], we propose a simple but effective multistage contrastive learning approach for point clouds in an unsupervised setting. As shown in Fig. 3, given the augmented point clouds input $P_{in,1}$ and $P_{in,2}$, we feed them into the backbone encoder. We add five additional MLP heads with region-level max-pooling to obtain region-level segmentation predictions at the $m_{th}$ backbone stage, denoted as $\textbf{f}_{g,m}$ and $\textbf{f}_{h,m}$, respectively (five stages in our case, denote $M$ as the total network stages. i.e., $M$ = 5). After we apply the MLP heads to the extracted features at different stages, we can obtain the hierarchical feature embeddings of the fine-tuned downstream 3D scene understanding network. Unlike existing pixel-level [124], or point-level [124, 125] for point cloud contrastive learning, our proposed contrastive learning performs at the region level. Therefore, it will help to obtain more holistic regional information. The obtained regional information will be complementary to the point-level information. The region-level semantic contrastive loss is formulated as:
$L^{m}_{contrast}=-\frac{1}{\textbf{S}_{p}}\sum_{(a,b)\in S_{p}}log\frac{H_{i,j}exp(\textbf{f}_{g,m}^{a}\cdot\textbf{f}_{h,m}^{+}/\tau)}{\sum_{(\cdot,c)\in S_{p}}H_{i,j}exp(\textbf{f}_{g,m}^{a}\cdot\textbf{f}_{h,m}^{-}/\tau))},$ (11)
where $(a,b)\in\textbf{S}_{p}$ are latent confident predicted positive region pairs, and $(a,c)\in\textbf{S}_{p}$ are latent negative region pairs. As mentioned before, $H_{i,j}$ is designed to eliminate contrastive learning candidates with low confidence degrees. Reliable region-level contrastive learning is only applied to confident predictions by the network. After applying contrastive losses for instance discrimination, semantically similar features with high confidence levels in the latent space will become more adjacent, whilst semantically distinct ones will be separated, which will be shown in Fig. 13. Note that although a recent work GPC [121] has proposed methods to perform contrastive learning on the point clouds in a SSL manner, our work is different from their method in two aspects. Firstly, our contrastive learning is conducted at region level while GPC conducts contrastive learning at the point level. Secondly, GPC focuses on the selection of the positive and negative point-set samples to perform contrastive learning in a pseudo-label supervised manner on two different 3D scene samples, while we focus on the unsupervised contrastive learning which disentangles different feature representations in the latent space on the same augmented 3D scene sample, guided by confident network predictions, and it learns hirachical regional object information.
The final proposed multi-stage contrastive learning loss is formulated as the sum of losses at every network stage:
$L_{multi}=\sum_{m=1}^{M}L_{contrast}^{m},\vspace{-0.1mm}$ (12)
After applying the multi-stage contrastive loss $L_{multi}$, the output at each stage of the network will provide more distinctive representations to attain better performance. From our ablation experiments, the performances can be boosted by applying multi-stage contrastive loss. We apply the typical self-training-based approach, where the high-confidence predictions are retained as the pseudo label for the network training in the subsequent iteration. We adopt the typical cross-entropy loss for the self-training which is denoted as $L_{ST}$. Considering $E_{sum}$ (see Eq. 8) for the two augmented scenes $P_{in,1}$, $P_{in,2}$, and $L_{multi}$ (see Eq. 12), we formulate the overall loss $L_{unlabeled}$ for the WS3D training with unlabeled data: $L_{unlabeled}=L_{ST}+\alpha E_{sum,2}+\beta L_{multi}$.
3.2.3 Supervised Learning for labeled data
We also guide network optimizations by using supervision from the labeled data. As shown in Fig. 3, we use the cross-entropy loss $L_{ce}$ to guide the supervised learning on the labeled data in the supervised branch. The loss term for the WS3D training with the labeled data is $L_{labeled}$. For the task of semantic segmentation, the loss for the labeled data $L_{labeled}$ includes the cross-entropy loss. For the backbone network selection, we follow previous work with the sparse convolutional network design for semantic [59] and instance segmentation [90], respectively. Our detection network is designed based on widely adopted VoteNet [126, 99]. We also use Dice [127] loss to guarantee tighter aggregations of points within the same cluster, and strict geometric separations of points in diverse clusters [99].
3.2.4 The Overall Optimization Loss Function
Leveraging our proposed region-level energy-based loss and region-level contrastive learning, the network can make maximum use of the unlabeled data for better feature learning to boost performance in the fine-tuning stage. As shown in Fig. 3, for semantic segmentation and instance segmentation, we fine-tune the network in an end-to-end manner for both supervised and unsupervised branches to make full use of labeled and unlabeled data. The weight value of the unsupervised loss is denoted as $\lambda_{u}$. The overall optimization function $L_{total}$ is formulated as follows, and we set $\lambda_{u}=2$ empirically:
$L_{total}=L_{labeled}+\lambda_{u}L_{unlabeled}.$ (13)
Figure 6: The captured 2D and 3D region proposals. It is demonstrated that more object proposals can be captured by our proposed combined unitive 2D/3D proposal generation approach.
Figure 7: The final ScanNet [20] object detection results. Differences in predictions are highlighted in the rectangles.
Figure 8: The object detection comparisons on KITTI [128] validation set. It can be demonstrated that our proposed WS3D++ provides very accurate bounding box predictions qualitatively compared with the previous state-of-the-art merely using 2D bounding boxes, which demonstrates the effectiveness of using 3D regional as well as 3D object-level information in facilitating effective feature representation learning.
Figure 9: Segmentation result comparisons with CLIP prompts for the indoor ScanNet benchmark. It can be seen from our results that clear object-level information can be captured with the designed visual prompts, which demonstrates the effectiveness of our designed hierarchical feature-aligned representation learning approaches.

### v2

@@ 3.1 Hierarchical Vision-Language Knowledge Associations and Distillations for Pre-training
Multi-view rendering. To obtain paired 2D-3D representations, we propose leveraging multi-view rendering to obtain paired 2D views from 3D point cloud scenes. The pairing process consists of two steps, the first is to convert point cloud scenes into meshes and the second is to render 2D images based on the different views of the 3D meshes. In terms of point-to-mesh transformation, we utilize the Delaunay triangulation approach [93, 94] to convert the point cloud into meshes, which is demonstrated as a very effective method for surface reconstruction. It connects the points in point cloud scenes by forming triangles that satisfy the Delaunay criterion which guarantees no point lies inside the circumcircle of any triangle. This method generates a triangle mesh that approximates the surface of the point cloud [95]. In terms of mesh-to-image transformation, we leverage the rendering pipeline including the vertex transformation, projection, and rasterization, as well as shading. We directly use the rendering library OpenGL [96] to render images from meshes. The process involves projecting the 3D vertices onto a 2D image plane based on camera parameters and applying shading and lighting calculations of 3D meshes to determine the specific color of each pixel. By our simple rendering design, the world-to-camera extrinsic transformation matrix $T_{e}$ containing both rotation and translation information between the 2D pixels and 3D points can be easily obtained.
@@
Figure 8: The object detection comparisons on KITTI [98] validation set. It can be demonstrated that our proposed WS3D++ can provide very accurate bounding box predictions qualitatively compared with the previous state-of-the-art merely using 2D bounding boxes, which demonstrates the effectiveness of using 3D regional as well as 3D object-level information in facilitating effective feature representation learning.
Figure 9: Segmentation result comparisons with CLIP prompts for the indoor ScanNet benchmark. It can be seen from our results that clear object-level vision-language matched information can be captured with the designed visual prompts. It reveals the effectiveness of our designed hierarchical visual linguistic feature-aligned representation learning approaches.
@@
United 2D and 3D Proposal Generation. As shown in Fig. 6, according to our experiments, we found that it is difficult to directly find the object-level information based on the 2D proposals merely due to the large information loss while rendering, the proposals provided by the 2D region proposal network (RPN) [99] can not effectively capture the 3D object information within the holistic scene. On the other hand, merely relying on the 3D proposals provided by the 3D RPN [100] still can not guarantee accurate proposal generation for the fact that some objects are too adjacent in their geometry. To this end, we propose to leverage the union of 2D and 3D RPN to capture intact and all-inclusive holistic proposals within the 3D scenes. Denote the region proposals as $R_{2D}$ and $R_{3D}$ respectively, the final holistic proposal generation $R_{H}$ is formulated as $R_{H}=R_{2D}\cup R_{3D}$. According to our experiments, this simple design can considerably boost the performance for the fact that the objects are more clustered and closely distributed within the indoor 3D scenes. Taking the union of 2D and 3D object proposals into consideration also guarantees that the optimization merely considers the regions where the object really exists and prevents the models from taking the pure background into consideration during the optimizations. According to our previous experiments, although the average precision and recall can merely achieve 38.7/45.9% and 47.6/56.8% for 2D and 3D object proposal generation, respectively, the 3D scene parsing performance can still be well-maintained. By combining the 2D proposal with the 3D using the union operation while omitting the duplicated 2D proposals, the object proposal generation precision and recall can achieve a very high score of 65.2% and 76.8%, respectively, which demonstrates the superior effectiveness of them in generating region proposals.
In the next stage, we perform vision-language matched contrastive candidate selection at both the object level and the scene level. The object-level vision-language matching makes alignment between the individual object and the descriptive word. While the scene-level vision-language matching make alignment between the whole scene with the descriptive sentence. In this way, the vision-language matched candidate can be selected for effective constrastive optimizations taking both global and local feature representation into account.
Although the accuracy of proposal generation is not that high, our proposed approach can guarantee the final scene parsing performance as reported and demonstrated in Tables 1-8. The reason behind the phenomenon can be explained by that as long as the region proposal can be found, our framework can leverage the well-aligned representation in the vision language model (CLIP) to boost the final 3D semantic scene parsing performance, both in the open-vocabulary and closed-set scenarios. As validated by the previous experimental results and our qualitative validation in Figure 6, the quality of object proposal can be guaranteed, if the domain gap is not that large.
Global 2D Scene-to-Sentence Matching. In the first place, we perform the global scene-to-sentence matching. For the rendered 2D image of the 3D scene, we utilize the GPT-4 to obtain the direct sentence-level description of the holistic scene. The sentence is given by the most similar description generated by GPT-4. Benefiting from the previous effective multi-view rendering designs, our proposed approach can have a holistic multi-view abstraction of the 3D environment. Denote the extract global scene-level visual feature as $V_{global}$ and denote the global text-level feature as $L_{global}$, then we can evaluate the alignment between visual and linguistic feature utilizing the similarity between $V_{global}$ and $L_{global}$.
Local 2D Object-to-Word Matching. At the next step, we perform local object-to-word matching. In contrast to the global scene to scentence matching, here we conduct local object-to-word matching, which establishes and enhances vision-language association at more fine-grained object level and word level. The association is established at the object level for visual part and at the word level for language part. Specifically, we use the CLIP image encoder to generate a set of local visual feature embeddings $V_{l}ocal$ and utilize the text encoder to output a set of local word-level feature embeddings $L_{l}ocal$. In the next step, we calculate aggregated similarity using the cosine similarity $\mathcal{G}_{sim}$ between the 3D visually encoded features and the textual features:
$\textbf{S}(\textbf{V},\textbf{L})=\mathcal{G}_{sim}(V^{2D}_{gobal},L^{2D}_{gobal})+\mathcal{G}_{sim}(V^{2D}_{local},L^{2D}_{local})$ (1)
@@
2D-3D Visual Feature Distillation The main purpose is to conduct 2D-3D Visual Feature Distillation to obtain the aligned 2D-3D visual feature in the 2D-3D-language co-embedded feature space. We use the global 3D backbone to extract the 3D visual feature representation $V^{3D}_{global}$, and then utilize the $\mathcal{KL}$ divergence as the distillation loss to further distill the vision-language aligned informative knowledge from 2D feature space to 3D. This process can also be interpreted as the 2D-3D explicit feature alignment/distillation process. Compared with the mean square error loss such as the $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ losses, the $\mathcal{KL}$ divergence has improved regression capacity and ensured smoother gradient, which to some extent overcomes overfitting problems while distilling knowledge. The $\mathcal{KL}$ distillation loss $\mathcal{L}^{Dist}_{\mathcal{KL}}$ finally operates on the final two normalized 2D/3D vector for feature alignment:
@@
Contrastive Language-Vision Optimizations. Note that when conducting contrastive learning, we regard textual features as anchors because textual descriptions are highly semantic and contain rich information, whereas the images contain too much low-level information and pixel-level details. Denote the $\textbf{V}^{3D,+}_{global}$ and the $\textbf{L}^{3D,-}_{global}$ as the positive and the negative feature with respect to the anchor textual feature $\bm{F}_{T}^{a}$, respectively, the designed contrastive discrimination loss is formulated as follows:
@@
According to our extensive experiments, our simple pre-training approach provides the well-aligned vision-language-3D aggregated co-embedding. It considerably facilitates vision-language-associated knowledge transfers from 2D to 3D, boosting both the final label efficiency and the final recognition capacity of unseen novel categories.
@@ 3.2 Region-Aware Fine-tuning
During the fine-tuning stage, our proposed framework consists of three subparts for the network optimization: 1. Unsupervised energy-based loss guided by boundary awareness and highly confident network predictions for unlabeled data, which is discussed in our original ECCV conference work [4]; 2. Unsupervised multi-stage region-level contrastive learning with highly confident predictions for unlabeled data. 3. Supervised semantic contrastive learning for labeled data. The three important modules above are integrated jointly into the optimization function for network training to accomplish the final downstream detection or segmentation tasks with a very limited labeled data, with all remaining data unlabeled.
Figure 10: The segmentation result visualizations and comparisons with CLIP text encoder prompts for the outdoor KITTI benchmark. It can be demonstrated the final foreground object awareness can be clearly captured as compared with the previous SOTAs approach PLA [69]. Meanwhile, as shown in the last three columns, we can provide clear segmentation for the corresponding visual objects based on the textual prompts. The results further demonstrate that vision-language aligned representations can be effectively and sufficiently learnt.